OpenStudy (anonymous):

What is the value of the discriminant of 2x^2 – 3x – 7 = 0?

OpenStudy (anonymous):

9-4(2)(-7) = 9-(-56)=65

OpenStudy (anonymous):

the discriminant is the part under the radical in the quadratic so it would be b^2-4ac

OpenStudy (anonymous):

oh! awesome! thank you! :)

OpenStudy (anonymous):

use quadratic formula for this. :D

OpenStudy (amistre64):

i tend to get confused with discriminant and determinant. I think they were have a sale on d words that day ....

OpenStudy (anonymous):

i call it discriminant. xD

OpenStudy (anonymous):

Oh wow...I just realized I don't know the difference either...